OpenStudy (alexwee123):

help me with this SAT 2 question If three numbers log(sub a), log(sub b), and log(sub c) in that order form a arithmetic progression, which of the following is true? a) b=ac b) c=ab c)b=(a+c)/2 d)b^2=ac e)b^2=a+c

OpenStudy (anonymous):

from the property of arithmetic progression: 2*log(sub a)=log(sub a)+log(sub c) log[(sub b)^2]=log(sub a*sub c) compare it and find the relation between a,b,c it will be b^2=ac

OpenStudy (alexwee123):

makes total sense now thanks :D

OpenStudy (anonymous):

welcome